# Exploring the Energy Landscape of RBMs: Reciprocal Space Insights into Bosons, Hierarchical Learning and Symmetry Breaking

### v1

@@ 5.2 Hierarchical Learning
Hierarchical learning is the mechanism by which different pieces of deep learning models learn different levels of abstraction. In the case of deep neural networks, it has been shown that deeper layers in the network learns more complex features [32]. In diffusion models, it has been shown that the feature hierarchies occur at different timescale during the denoising process [28]. In GANs it has been shown that the class typically clusters around a subset of eigenvectors in the latent space [33]. In the case of RBM, it has been shown that during training, the eigenvectors align with the principal components of the dataset [27], which we say is similar to hierarchical learning. We can probe this claim by rotating the eigenvector matrix $U$. To do so, notice that when sampling from an RBM, the last step consists on sampling from $\sigma(\bm{W}\bm{h}+\bm{b})=\sigma(\bm{U}\bm{\Sigma}\bm{V}^{t}\bm{h}+\bm{b})$. Consider a rotation matrix $\mathcal{R}$, and define the rotated orthogonal matrix $\bm{U}_{R}=\mathcal{R}\bm{U}\mathcal{R}^{t}$. We build a rotation matrix using the method described in Ref. [31]. This method performs rotations in $n-$dimensions around any arbitrary $(n-2)$-dimension subspace. In Fig. 6 we show a sample generated by block Gibbs. We also show the effect on the samples of rotating by $\pi$ the eigenvector matrix $\bm{U}$ around all but the two dimensions associated with the largest singular values. Similarly, we show the case where we do 10 consecutive rotations around any $(N-2)$-dimension where the dimensions associated with the five largest singular values are always included and the remaining $N-7$ dimensions are randomly selected. Notice how rotating the eigenvectors associated with the largest singular values has a global effect on the image, whereas rotating eigenvectors associated with any but the principal eigenvectors has a local effect on the image. In this sense, the eigenvectors are associated with different feature hierarchies.
@@ 5.3 Symmetry breaking in energy landscape
which resembles the free energy in Landau theory. To validate this approximation, in Fig. 7 we plot the evolution of $\mu_{i}$ vs epochs for $i=1$ and $i=M+1$. We also include the projection of Gibbs sampled data and test data projected onto the energy landscape, which we can assume correspond to the effective potential minimum. For $i=1$ we observe a symmetry breaking occurring at early epochs where the free energy evolves from an harmonic well to a double well, reminiscent of a ferromagnetic phase transition. The fact that $\mu_{i}$ is located on opposite side of the local minimum with respect to the origin indicate that the double well is non-symmetric. Hence the interplay between the energy function and the constraint potential during training induces symmetry breaking in reciprocal space. Specifically, an increase in magnitude of the prefactor $-\lambda_{i}/2$ together with a shift of value of the minimum position of the constraint potential, $\mu_{i}$, from zero to non-zero leads to the symmetry breaking. In contrast, for $i=M+1$ we observe that the corresponding mode becomes relatively flat.
This characteristic persists across different hidden layer sizes, as illustrated in Fig. 8. As the hidden layer size increases, the constraint potential minimum, $\mu_{i}$, shifts closer to the effective potential minimum obtained by the test set and block Gibbs sampling. In general, the symmetry breaking is explicit, since the free energy is asymmetric. In the special case where the number of hidden nodes equals the number of visible nodes, the constraint potential minimum tends to align closely with the energy function saddle point. However, the free energy is still asymmetric, in general, due to the first- and third-order terms in the free energy.
@@ 7 Acknowledgements
JQTM is thankful to Jorge Fernandez de Cossio, Daniel Miravet, Mehdi Drissi and Ejaaz Merali for useful discussions. JQTM acknowledges a Mitacs Elevate Postdoctoral Fellowship (IT39533) with Perimeter Institute for Theoretical Physics. We gratefully acknowledge funding from the National Research Council (Canada) via Agreement AQC-002. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research, Innovation and Science. The University of Virginia acknowledges support from NSF 2212550 OAC Core: Smart Surrogates for High Performance Scientific Simulations and DE-SC0023452: FAIR Surrogate Benchmarks Supporting AI and Simulation Research. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).

### v2

@@ 5.2 Hierarchical Learning
Hierarchical learning is the mechanism by which different pieces of deep learning models learn different levels of abstraction. In the case of deep neural networks, it has been shown that deeper layers in the network learns more complex features [44]. In diffusion models, it has been shown that the feature hierarchies occur at different timescale during the denoising process [40]. In GANs it has been shown that the class typically clusters around a subset of eigenvectors in the latent space [45]. In the case of RBM, it has been shown that during training, the eigenvectors align with the principal components of the dataset [37], which we say is similar to hierarchical learning. We can probe this claim by rotating the eigenvector matrix $U$. To do so, notice that when sampling from an RBM, the last step consists on sampling from $\sigma(\bm{W}\bm{h}+\bm{b})=\sigma(\bm{U}\bm{\Sigma}\bm{V}^{t}\bm{h}+\bm{b})$. Consider a rotation matrix $\mathcal{R}$, and define the rotated orthogonal matrix $\bm{U}_{R}=\mathcal{R}\bm{U}\mathcal{R}^{t}$. We build a rotation matrix using the method described in Ref. [43]. This method performs rotations in $n$-dimensions around any arbitrary (n2)(n-2)-dimension subspace. In Fig. 6 we show a sample generated by block Gibbs. We also show the effect on the samples of rotating by $\pi$ the eigenvector matrix $\bm{U}$ around all but the two dimensions associated with the largest singular values. Similarly, we show the case where we do 10 consecutive rotations around any (N2)(N-2)-dimension where the dimensions associated with the five largest singular values are always included and the remaining N7N-7 dimensions are randomly selected. Notice how rotating the eigenvectors associated with the largest singular values has a global effect on the image, whereas rotating eigenvectors associated with any but the principal eigenvectors has a local effect on the image. In this sense, the eigenvectors are associated with different feature hierarchies.
@@ 5.3 Symmetry breaking in energy landscape
which resembles the free energy in Landau theory. To validate this approximation, in Fig. 7 we plot the evolution of $\mu_{i}$ vs epochs for $i=1$ and $i=M+1$. We also include the projection of Gibbs sampled data and test data projected onto the energy landscape, which we can assume correspond to the effective potential minimum. For $i=1$ we observe a symmetry breaking occurring at early epochs where the free energy evolves from an harmonic well to a double well, reminiscent of a ferromagnetic phase transition. The fact that $\mu_{i}$ is located on opposite side of the local minimum with respect to the origin indicate that the double well is non-symmetric. Hence the interplay between the energy function and the constraint potential during training induces symmetry breaking in reciprocal space. Specifically, an increase in magnitude of the prefactor $\lambda_{i}/2$ together with a shift of value of the minimum position of the constraint potential, $\mu_{i}$, from zero to non-zero leads to the symmetry breaking. In contrast, for $i=M+1$ we observe that the corresponding mode becomes relatively flat.
This characteristic persists across different hidden layer sizes, as illustrated in Fig. 8. As the hidden layer size increases, the constraint potential minimum, $\mu_{i}$, shifts closer to the effective potential minimum obtained by the test set and block Gibbs sampling. In general, the symmetry breaking is explicit, since the free energy is asymmetric. In the special case where the number of hidden nodes equals the number of visible nodes, the constraint potential minimum tends to align closely with the energy function saddle point. However, the free energy is still asymmetric, in general, due to the first- and third-order terms in the free energy. This result is in agreement with Ref. [34], where the authors also identified a mechanism of symmetry breaking by studying the equilibrium properties of RBMs, and further showed the transition from a paramagnetic phase to a ferromagnetic phase driven by the learning dynamics.
@@
6 Discussion
In the preceding sections we established a clear connection between RBMs and diffusion process, but a direct link with diffusion models is pending. Such a connection could have a great impact in Quantum Machine Learning, as it could enable the use of quantum annealers to implement diffusion models with faster sampling. The main obstacle is that diffusion models, in general, do not satisfy the detailed balance condition. Nonetheless, different approaches have been proposed. For instance, in Ref. [46] the author proposes an adiabatic diffusion model, while in Ref. [47] the authors introduce consistency models whereby a separate neural network is trained to mimic the diffusion model, enabling one-step sample. In addition, we established a clear and novel connection between RBMs and coupled Bosons, which may provide new pathways that relate RBM with problems in many-body physics. We showed that sampling the weight matrix from a Gaussian distribution renders the RBM weight distribution rotationally symmetric. For an infinite size hidden layer, the reciprocal variables associated with the non-zero singular values are Gaussian distributed. In this limit, the associated Bosonic system decouples into a set of uncoupled harmonic oscillators, with eigenvalues $E_{i}=4\pm\lambda_{i}\beta$. There is a critical singular value, $\lambda_{c}=4/\beta$, for when $\lambda>\lambda_{c}$, the bound states hinder the diffusion process from reaching the Boltzmann distribution. Consequently, RBM configuration can become trapped in non-Boltzmann states, leading to biased gradient estimates during maximum-likelihood training. It is possible, however, that after a few training steps, once this symmetry is broken the reciprocal variables are no longer Gaussian distributed and the coupled Bosons bound states disappear. Testing this is left for future work.
Our results also provide a theoretical understanding for known rules-of-thumb in RBM which renders the RBM rotational symmetry and provided:
  • RBM layer sizes and weights variance. The size of the hidden layer typically is less or equal to the visible layer. Our results show that for a large difference between hidden and visible layer sizes and/or weights with large variance, the energy landscape can be unstable at initialization.
  • Instabilities in Gaussian RBMs. For Gaussian‐Gaussian RBMs, the energy function is effectively a set of hyperbolic paraboloids in reciprocal space, while the corresponding constraint potential is relatively flat. This allows the reciprocal variables to drift, rendering Gaussian RBMs especially prone to instability and difficult to train.
  • Regularizers. The use of regularizers during training is a common approach when training generative models. Particularly, the L2-norm is used when training an RBM. In maximum posterior training, one typically maximizes
    $\log P(\text{weights}|\text{data})=\log P(\text{data}|\text{weights})+\log P(\text{weights})-P(\text{data})\;.$ (57)
    where the term $\log P(weights)$ is the L2-norm, and can be understood as a constraint on the model enforcing rotational symmetry in the RBM. However, we note that breaking this symmetry is important for hierarchical learning, but it does not guarantee that the model captures the data set distribution. In some cases, the symmetry may break without improving performance. A more tailored L2-norm, where the constraints are applied with different weights on linearly independent directions of the weight matrix can encourage symmetry breaking in a way that enhances the learning.
Rotational symmetry in RBMs can be viewed as analogous to starting a sculpture from a perfectly symmetrical block of material. Training corresponds to carving away symmetries until the model captures the desired data structure. This underscores the relevance of rotational symmetry in RBMs. However, as we stressed before, under certain conditions this symmetry may hinder the RBM training. In Ref. [22] the authors show a symmetry breaking from a phase the random guess phase where there is no symmetry breaking and the data do not provide any useful information to bias the weight’s direction to a phase where the reverse symmetry breaks. In Ref. [23] the authors show that RBMs undergo a second order phase transition during training at early epochs. It remains to be unvelied the full phase diagram in RBMs. We leave this exciting experiment for future work, as this will require intensive numerical results.
@@ 8 Acknowledgements
JQTM is thankful to Jorge Fernandez de Cossio, Daniel Miravet, Mehdi Drissi and Ejaaz Merali for useful discussions. The authors thank the reviewers for their comments and discussion. JQTM acknowledges a Mitacs Elevate Postdoctoral Fellowship (IT39533) with Perimeter Institute for Theoretical Physics. We gratefully acknowledge funding from the National Research Council (Canada) via Agreement AQC-002. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research, Innovation and Science. The University of Virginia acknowledges support from NSF 2212550 OAC Core: Smart Surrogates for High Performance Scientific Simulations and DE-SC0023452: FAIR Surrogate Benchmarks Supporting AI and Simulation Research. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).